Your starship, the Aimless Wanderer, lands on the mysterious planet Mongo. As chief scientist-engineer, you make the following measurements: A 2.50-kg stone thrown upward from the ground at 11.5 m/s returns to the ground in 8.08 s; the circumference of Mongo at the equator is 2.00 x 105 km; and there is no appreciable atmosphere on Mongo. The starship commander, Captain Confusion, asks for the following information: (a) What is the mass of Mongo? x kg Enter a number (b) If the Aimless Wanderer goes into a circular orbit 30,600 km above the surface of Mongo, how many hours will it take the ship to complete one orbit?
Answer #1

using 2nd equation of motion along vertical

h = ut - 0.5 gt^2

0 = 11.5* 8.08 - 0.5* g *8.08^2

g = 2.847 m/s^2

now,

C = 2 pi R

2*10^8 = 2* 3.14* R

R = 3.185*10^7 m

Acceleration due to gravity can be determined by

G M / R^2 = 2.847

6.67*10^-11* M / (3.185*10^7) ^2 = 2.847

M = 4.329*10^25 kg

==========

b)

orbital velocity,

v = sqrt ( GM / ( R + h))

v = sqrt ( 6.697*10^-11* 4.329*10^25 / ( 3.185*10^7 + 30600000) )

v = 6800 m/s

Time taken by aimless wanderer

t = 2 pi ( R + h) /v

t = 2*3.14* ( 3.185*10^7 + 30600000) / 6800

t = 57675.88 seconds

t = 16.02 hours

=======

Comment in case any doubt, will reply for sure. Goodluck